In a classroom demonstration, the pressure inside a soft drink can is suddenly reduced to essentially zero. Assuming the can to be a cylinder with a height of 10 cm and a diameter of 6.2 cm , find the net inward force exerted on the vertical sides of the can due to atmospheric pressure.
Answer #1

Given,

height of cylinder, h = 10 cm = 0.10 m

diameter of the cylinder, d = 6.2 cm = 0.062 m

to find, the net inward force, F = ?

Now,

Pressure (P) = Force (F) / Area (A)

F = P*A

Here, P = 1x10^5 (Atmospheric pressure) and
A = Lateral surface Area = pi*d*h
   = pi*0.062*0.10

   = 0.01948

Force, F = 1x10^5 x (0.01948)

F = 1948 N Answer